# Selective Classification Under Distribution Shifts

### v1

@@ 3.3 Analysis of rejection patterns
We use the same setup as in the numerical experiment in Section 3.1 (see also Fig. 2), and plot in Fig. 3 (a-1) the RC curves induced by the various confidence-score functions [^2] For the classifier considered, $RL_{\text{geo-M}}$ and $RL_{\text{conf-M}}$ have the same SC performance as $\|w\|_{2}=1$.. It is clear that $RL_{\text{geo-M}}$ performs the best. To better understand the difference between $RL_{\text{geo-M}}$ and other score functions, we study their rejection patterns: we visualize in Fig. 3 (b-1)&(c-1) the samples rejected at $0.8$ coverage for $RL_{\text{geo-M}}$ and $SR_{\text{max}}$, respectively; see visualization of other score functions in Appendix D, whose rejection patterns are similar to that of $SR_{\text{max}}$. An iconic feature of $RL_{\text{geo-M}}$ is that it prioritizes rejecting samples closer to decision boundaries, whereas SR-based scores prioritize rejecting samples close to the origin. Conceptually, the former rejection pattern is favorable, as the goal of SC is exactly to reject uncertain samples on which classifier’s decisions can be shaky. More precisely, the difference in rejection patterns implies at least two things: (1) $RL_{\text{geo-M}}$ could be advantageous when most classification errors occur near the decision boundaries; (2) $RL_{\text{geo-M}}$ may be superior even when test samples have a moderate level of distribution shifts with respect to training. For example, when the test set has a slightly different $\mathcal{D}_{\mathcal{X}|\mathcal{Y}}$ than the training set (see Cases 2 & 3 below), mistaken samples due to the shift tend to be close to the decision boundaries and thus can be successfully rejected. Fig. 3 (d-1) plots the histograms of the robustness radii (i.e., the $\ell_{2}$ distance of a sample to the closest decision boundary) of selected samples at $0.8$ coverage, where the robustness radius quantitatively captures the extent of $\mathcal{D}_{\mathcal{X}|\mathcal{Y}}$ shift SC can tolerate: while the selected samples using $RL_{\text{geo-M}}$ uniformly have nonzero robust radii, all other score functions lead to zero robustness radii for the worst samples, implying sensitivity to $\mathcal{D}_{\mathcal{X}|\mathcal{Y}}$ shifts. [^3] The intuition on why our notions of margins work for Type B errors is different: there since $\mathbf{x}$ assumes a label outside the known set, we expect no clear winner in the raw logits.
@@ Acknowledgments
Liang H. and Sun J. are partially supported by NIH fund R01NS131314. Peng L. and Sun J. are partially supported by NIH fund R01CA287413. The authors acknowledge the Minnesota Supercomputing Institute (MSI) at the University of Minnesota for providing resources that contributed to the research results reported in this article. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ Appendix B Asymptotic behaviors of $SR_{\text{max}}$, $SR_{\text{doctor}}$, and $SR_{\text{ent}}$
$\log(-SR_{\text{ent}}(\lambda\mathbf{z}))$ $\sim-\log\sum_{j}e^{\lambda(z^{(j)}-z^{(1)})}+\lambda(z^{(2)}-z^{(1)})+\log(-\lambda(z^{(2)}-z^{(1)}))$ (39)
$\sim-\log(1+e^{\lambda(z^{(2)}-z^{(1)})})+\lambda(z^{(2)}-z^{(1)})\quad\quad(\text{$(\log(1+x)\sim x$ as $x\to 0$})$ (40)
@@
Table 5: Evalutaion of $s_{1}$ and $s_{2}$ using popular OOD metrics. The better numbers are highlighted in bold.
OOD metric $s_{1}$ $s_{2}$
AUROC ($\uparrow$) 0.765 0.944
AUPR ($\uparrow$) 0.987 0.997
FPR@TPR=0.95 ($\downarrow$) 0.816 0.279

### v2

@@ 3.3 Analysis of rejection patterns
We use the same setup as in the numerical experiment in Section 3.1 (see also Fig. 2), and plot in Fig. 3 (a-1) the RC curves induced by the various confidence-score functions [^2] For the classifier consideblue, $RL_{\text{geo-M}}$ and $RL_{\text{conf-M}}$ have the same SC performance as $\|w\|_{2}=1$.. It is clear that $RL_{\text{geo-M}}$ performs the best. To better understand the difference between $RL_{\text{geo-M}}$ and other score functions, we study their rejection patterns: we visualize in Fig. 3 (b-1)&(c-1) the samples rejected at $0.8$ coverage for $RL_{\text{geo-M}}$ and $SR_{\text{max}}$, respectively; see visualization of other score functions in Appendix D, whose rejection patterns are similar to that of $SR_{\text{max}}$. An iconic feature of $RL_{\text{geo-M}}$ is that it prioritizes rejecting samples closer to decision boundaries, whereas SR-based scores prioritize rejecting samples close to the origin. Conceptually, the former rejection pattern is favorable, as the goal of SC is exactly to reject uncertain samples on which classifier’s decisions can be shaky. More precisely, the difference in rejection patterns implies at least two things: (1) $RL_{\text{geo-M}}$ could be advantageous when most classification errors occur near the decision boundaries; (2) $RL_{\text{geo-M}}$ may be superior even when test samples have a moderate level of distribution shifts with respect to training. For example, when the test set has a slightly different $\mathcal{D}_{\mathcal{X}|\mathcal{Y}}$ than the training set (see Cases 2 & 3 below), mistaken samples due to the shift tend to be close to the decision boundaries and thus can be successfully rejected. Fig. 3 (d-1) plots the histograms of the robustness radii (i.e., the $\ell_{2}$ distance of a sample to the closest decision boundary) of selected samples at $0.8$ coverage, where the robustness radius quantitatively captures the extent of $\mathcal{D}_{\mathcal{X}|\mathcal{Y}}$ shift SC can tolerate: while the selected samples using $RL_{\text{geo-M}}$ uniformly have nonzero robust radii, all other score functions lead to zero robustness radii for the worst samples, implying sensitivity to $\mathcal{D}_{\mathcal{X}|\mathcal{Y}}$ shifts. [^3] The intuition on why our notions of margins work for Type B errors is different: there since $\mathbf{x}$ assumes a label outside the known set, we expect no clear winner in the raw logits.
@@ Acknowledgments
Liang H. and Sun J. are partially supported by NIH fund R01NS131314. Peng L. and Sun J. are partially supported by NIH fund R01CA287413. The authors acknowledge the Minnesota Supercomputing Institute (MSI) at the University of Minnesota for providing resources that contributed to the research results reported in this article. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. This research is also part of AI-CLIMATE: “AI Institute for Climate-Land Interactions, Mitigation, Adaptation, Tradeoffs and Economy,” and is supported by USDA National Institute of Food and Agriculture (NIFA) and the National Science Foundation (NSF) National AI Research Institutes Competitive Award no. 2023-67021-39829.
@@ Appendix B Asymptotic behaviors of $SR_{\text{max}}$, $SR_{\text{doctor}}$, and $SR_{\text{ent}}$
$\log(-SR_{\text{ent}}(\lambda\mathbf{z}))$ $\sim-\log\sum_{j}e^{\lambda(z^{(j)}-z^{(1)})}+\lambda(z^{(2)}-z^{(1)})+\log(-\lambda(z^{(2)}-z^{(1)})).$ (39)
Let’s write $x\triangleq-(z^{(2)}-z^{(1)})$. The last two terms in Eq. 39 can be re-written as $-x+\log(x)$. Since $\lim_{x\to\infty}\frac{\log(x)}{x}=0$, we thus have $-x+\log(x)=-x(1+o(1))$ as $x\to\infty$, and hense $-x+\log(x)\sim-x$ by the definition of the asymptotic equivalence. Therefore, we have:
$\log(-SR_{\text{ent}}(\lambda\mathbf{z}))$ $\sim-\log(1+e^{\lambda(z^{(2)}-z^{(1)})})+\lambda(z^{(2)}-z^{(1)})$ (40)
@@
Table 5: Evaluation of $s_{1}$ and $s_{2}$ using popular OOD metrics. The better numbers are highlighted in bold.
OOD metric $s_{1}$ $s_{2}$
AUROC ($\uparrow$) 0.765 0.944
AUPR ($\uparrow$) 0.987 0.997
FPR@TPR=0.95 ($\downarrow$) 0.816 0.279
The commonly used evaluation metrics for OOD detection do not reflect the classification performance (Franc et al., 2023b). Here we provide a quantitative supporting example, in comparison with the RC curve for generalized SC.